) Answer the following questions: (10 points) Is it possible for one computer with a low MIPS rating to have a better performance in practice than a computer with a high MIPS rating? Explain a. What is the MIPS rating of this computer? Given that a program is run on a computer clock cycle time instructions with 1 cycle b. with the following parameters: 10 ns 70% 20% 1096 instructions with 2 cycles instructions with 3 cycles
Answer #1

a.

MIPS or Million Instructions Per Second is an obsolete method to measure a computer's speed and power. It roughly measures no. of instructions that a computer can execute per second. However, different instructions require more or less time than others, and MIPS is limited to only to CPU speed whereas real applications take into account other factors too like I/O speed, processor architecture etc.

For example, a computer rated at 100 MIPS may be able to computer certain functions faster than another computer rated at 120 MIPS.

A machine with a high MIPS rating, therefore, might not run a particular application any faster than a machine with a low MIPS rating.

b.

clock cycle time = 10ns = 100 MHz

from the given instructions, CPI(clocks per instruction) = ((1*70)+(2*20)+(3*10))/100 = (70+40+30)/100 = 1.4

MIPS rating = Clock Rate/(CPI*106) = (100*106)/1.4*106 = 71.43